OpenStudy (anonymous):

Which of the following is one of the factors of 3x2 + 45x + 150? Answer x – 10 5 All of the above x + 5

OpenStudy (anonymous):

x - 10

OpenStudy (anonymous):

\[3 x^2+45 x+150=3 (x+5) (x+10) \]Answer = x + 5